# Evolution 6.0: Evolving Robotic Capabilities Through Generative Design

### v2

@@ Abstract
We propose a new concept, Evolution 6.0, which represents the evolution of robotics driven by Generative AI. When a robot lacks the necessary tools to accomplish a task requested by a human, it autonomously designs the required instruments and learns how to use them to achieve the goal. Evolution 6.0 is an autonomous robotic system powered by Vision-Language Models (VLMs), Vision-Language Action (VLA) models, and Text-to-3D generative models for tool design and task execution. The system comprises two key modules: the Tool Generation Module, which fabricates task-specific tools from visual and textual data, and the Action Generation Module, which converts natural language instructions into robotic actions. It integrates QwenVLM for environmental understanding, OpenVLA for task execution, and Llama-Mesh for 3D tool generation. Evaluation results demonstrate a 90% success rate for tool generation with a 10-second inference time, and action generation achieving 83.5% in physical and visual generalization, 70% in motion generalization, and 37% in semantic generalization. Future improvements will focus on bimanual manipulation, expanded task capabilities, and enhanced environmental interpretation to improve real-world adaptability.
@@ I Introduction
Unlike Industry 6.0, which operates in predefined industrial settings, we propose Evolution 6.0 illustrated in Figure 1, a framework that extends these capabilities to unpredictable environments, such as Mars, where tool scarcity and environmental variability are major challenges. Evolution 6.0 empowers autonomous robots to analyze unfamiliar surroundings, interpret contextual cues via VLM, and generate actionable solutions with VLA models. This enables robots to autonomously determine, design, and fabricate tools as needed, eliminating dependency on predefined instructions and enhancing operational flexibility.
Figure 2: Evolution 6.0 system architecture
@@ II Related Work
III System Architecture
The Evolution 6.0 is illustrated in Figure 2. Our system integrates two connected modules that work together to enable flexible robotic manipulation. Tool generation module focuses on creating situation-specific tools, while the Action generation module determines how to perform each task. This design adapts its actions based on observations of the environment and instructions expressed in everyday language. The approach uses QwenVLM for interpreting surroundings, OpenVLA for determining task steps, and LlamMa-Mesh for shaping three-dimensional tools. Through this arrangement, robots can operate effectively in environments that require tailored solutions and adaptive behaviors.
@@ III-A Tool generation module
The Tool Generation Module is responsible for analyzing the robot’s environment and autonomously generating the necessary tools to accomplish a given task. It begins with the Robot Scene Input, where sensors such as cameras capture the robot’s environment to identify objects requiring interaction. The captured data is processed through a Qwen2-VL-2B-Instruct, where a Vision Encoder extracts relevant visual features, which are then interpreted by the QwenLM decoder to generate contextual textual descriptions. Based on the scene analysis, the system formulates a prompt, such as “Create a 3D model of a knife,” which is passed to an Autoregressive Language Model (LlamMa-Mesh). This model synthesizes a 3D representation of the required tool in mesh format, containing vertex (v) and face (f) data that define the tool’s geometry. The generated mesh proceeds to the Mesh Rendering phase, where it is visualized and validated to ensure it meets the task requirements. Once validated, the finalized mesh is converted into G-Code, enabling the tool to be fabricated using a 3D printer. It is important to note that the system scales the tool by a factor, depending on the size of the item that needs to be manipulated in the scene. Figure 3 shows several tool examples produced by the LlamMa-Mesh.
@@ III-B Action generation module
To facilitate real-time processing, the QwenVLM and LlamMa-Mesh models were optimized by reducing numerical precision to int8, while the fine-tuned Vision-Language Action (VLA) model was kept in its original precision. This optimization, in conjunction with the utilization of a high-performance NVIDIA RTX 4090 graphics processing unit and streamlined communication protocols, enables the system to operate at a frequency of 5 Hz. This approach strikes an effective balance between computational efficiency and responsiveness, allowing Evolution 6.0 to meet the operational requirements of complex manipulation tasks.
@@ IV Dataset Collection and Training Pipeline
The dataset for training was collected using a UR10 robotic arm equipped with a Logitech C920e HD 1080p camera to capture data for robotic manipulation tasks. The dataset comprises 20 episodes, evenly divided into two tasks: cutting a piece of cake and picking and placing the cake. Throughout the data collection process, the robot’s gripper position, orientation, and state were recorded, alongside synchronized high-definition visual feedback. Each episode includes language instructions such as “Cut one piece of cake” or “Pick one piece of cake and place on plate.” The collected data was formatted into the RLDS (Reinforcement learning dataset) structure, ensuring compatibility with OpenVLA for efficient organization of actions, images, and task instructions, making it suitable for imitation learning and task-specific action prediction. The collected dataset was used to fine-tune an OpenVLA-7b model using a parameter-efficient LoRA approach, applying rank-32 adapters to optimize memory usage while adjusting a minimal set of trainable weights. The training setup included a batch size of 16, a learning rate of $5\times 10^{-4}$, and 4000 gradient steps, with image augmentation disabled. Training was conducted on a single A100 GPU, with periodic checkpointing. Key metrics such as training loss, action accuracy, and L1 loss were monitored to ensure model performance. Additionally, smoothed metrics were employed for stable progress tracking.
@@ V-1 Phase One: Tool Generation
The average inference time for these interpretations was approximately 4 seconds, contingent upon the specific scene. However, the model encountered challenges when presented with scenes containing multiple identical target objects or objects belonging to the same class as the target. For instance, when tasked with distinguishing between screws and bolts, the model consistently misinterpreted the prompt as ”Create the 3D model of the screwdriver,” which is incompatible with the bolt. Despite this limitation, the autoregressive LM demonstrated a high success rate in generating tools based on these prompts, achieving successful outputs in 9 out of 10 instances, with an average generation time of 10 seconds, varying according to the complexity of the tool. Moreover, the model struggled with generating complex curves or fillets in its outputs. As illustrated in accompanying Figure 3, all generated tools predominantly exhibited either sharp edges or simple circular cross-sections. This highlights a significant area for improvement in the model’s ability to handle more intricate geometrical features in tool design. The results are summarized in Table 1.
@@ V-2 Phase Two: Action Generation
Experimental results (summarized in Figure 5) highlight the module’s proficiency in physical and visual generalization and its resilience under varying conditions. However, there is significant room for improvement in motion and semantic generalization. For instance, when presented with instructions such as ”Cut the banana” or ”Place the tomato on the plate,” the module exhibited inconsistent behavior. Examples of these scenarios are illustrated in Figure 4.

### v4

@@ Abstract
We propose a new concept, Evolution 6.0, which represents the evolution of robotics driven by Generative AI. When a robot lacks the necessary tools to accomplish a task requested by a human, it autonomously designs the required instruments and learns how to use them to achieve the goal. Evolution 6.0 is an autonomous robotic system powered by Vision-Language Models (VLMs), Vision-Language Action (VLA) models, and Text-to-3D generative models for tool design and task execution. The system comprises two key modules: the Tool Generation Module, which fabricates task-specific tools from visual and textual data, and the Action Generation Module, which converts natural language instructions into robotic actions. It integrates QwenVLM for environmental understanding, OpenVLA for task execution, and Llama-Mesh for 3D tool generation. Evaluation results demonstrate a 90% success rate for tool generation with a 10-second inference time and action generation achieving 83.5% in physical and visual generalization, 70% in motion generalization, and 37% in semantic generalization. Future improvements will focus on bimanual manipulation, expanded task capabilities, and enhanced environmental interpretation to improve real-world adaptability.
@@ I Introduction
Unlike Industry 6.0, which operates in predefined industrial settings, we propose Evolution 6.0 illustrated in Fig. 1, a framework that extends these capabilities to unpredictable environments, such as Mars, where tool scarcity and environmental variability are major challenges. Evolution 6.0 empowers autonomous robots to analyze unfamiliar surroundings, interpret contextual cues via VLM, and generate actionable solutions with VLA models. This enables robots to autonomously determine, design, and fabricate tools as needed, eliminating dependency on predefined instructions and enhancing operational flexibility.
Figure 2: Evolution 6.0 system architecture.
@@ II Related Work
III System Architecture of Evolution 6.0
@@ III-A Tool generation module
The Tool Generation Module is responsible for analyzing the robot’s environment and autonomously generating the necessary tools to accomplish a given task. It begins with the Robot Scene Input, where sensors such as cameras capture the robot’s environment to identify objects requiring interaction. The captured data is processed through a Qwen2-VL-2B-Instruct, where a Vision Encoder extracts relevant visual features, which are then interpreted by the QwenLM decoder to generate contextual textual descriptions. Based on the scene analysis, the system formulates a prompt, such as “Create a 3D model of a knife,” which is passed to an Autoregressive Language Model (LlamMa-Mesh). This model synthesizes a 3D representation of the required tool in mesh format, containing vertex (v) and face (f) data that define the tool’s geometry. The generated mesh proceeds to the Mesh Rendering phase, where it is visualized and validated to ensure it meets the task requirements. Once validated, the finalized mesh is converted into G-Code, enabling the tool to be fabricated using a 3D printer. It is important to note that the system scales the tool by a factor, depending on the size of the item that needs to be manipulated in the scene. Fig. 3 shows several tool examples produced by the LlamMa-Mesh.
The Evolution 6.0 is illustrated in Fig. 2. Our system integrates two connected modules that work together to enable flexible robotic manipulation. Tool generation module focuses on creating situation-specific tools, while the Action generation module determines how to perform each task. This design adapts its actions based on observations of the environment and instructions expressed in everyday language. The approach uses QwenVLM for interpreting surroundings, OpenVLA for determining task steps, and LlamMa-Mesh for shaping three-dimensional tools. Through this arrangement, robots can operate effectively in environments that require tailored solutions and adaptive behaviors.
@@ III-B Action generation module
To facilitate real-time processing, the QwenVLM and LlamMa-Mesh models were optimized by reducing numerical precision to int8, while the fine-tuned Vision-Language Action (VLA) model was kept in its original precision. This optimization, in conjunction with the utilization of a high-performance NVIDIA RTX 4090 24Gb GPU and streamlined communication protocols, enables the system to operate at a frequency of 5 Hz. This approach strikes an effective balance between computational efficiency and responsiveness, allowing Evolution 6.0 to meet the operational requirements of complex manipulation tasks.
@@ IV Dataset Collection and Training Pipeline
The dataset for training was collected using a UR10 robotic arm equipped with a Logitech C920e HD 1080p camera to capture data for robotic manipulation tasks. The dataset comprises 20 episodes, evenly divided into two tasks: cutting a piece of cake and picking and placing the cake. Throughout the data collection process, the robot’s gripper position, orientation, and state were recorded, alongside synchronized high-definition visual feedback. Each episode includes language instructions such as “Cut one piece of cake” or “Pick one piece of cake and place on plate”. The collected data was formatted into the RLDS (Reinforcement learning dataset) structure, ensuring compatibility with OpenVLA for efficient organization of actions, images, and task instructions, making it suitable for imitation learning and task-specific action prediction. The collected dataset was used to fine-tune an OpenVLA-7b model using a parameter-efficient LoRA approach, applying rank-32 adapters to optimize memory usage while adjusting a minimal set of trainable weights. The training setup included a batch size of 16, a learning rate of $5\times 10^{-4}$, and 4000 gradient steps, with image augmentation disabled. Training was conducted on a single A100 GPU, with periodic checkpointing. Key metrics such as training loss, action accuracy, and L1 loss were monitored to ensure model performance. Additionally, smoothed metrics were employed for stable progress tracking.
@@ V-1 Phase One: Tool Generation
The average inference time for these interpretations was approximately 4 seconds, contingent upon the specific scene. However, the model encountered challenges when presented with scenes containing multiple identical target objects or objects belonging to the same class as the target. For instance, when tasked with distinguishing between screws and bolts, the model consistently misinterpreted the prompt as ”Create the 3D model of the screwdriver,” which is incompatible with the bolt. Despite this limitation, the autoregressive LM demonstrated a high success rate in generating tools based on these prompts, achieving successful outputs in 9 out of 10 instances, with an average generation time of 10 seconds, varying according to the complexity of the tool. Moreover, the model struggled with generating complex curves or fillets in its outputs. As illustrated in accompanying Fig. 3, all generated tools predominantly exhibited either sharp edges or simple circular cross-sections. This highlights a significant area for improvement in the model’s ability to handle more intricate geometrical features in tool design. The results are summarized in Table 1.
@@ V-2 Phase Two: Action Generation
Experimental results (summarized in Fig. 5) highlight the module’s proficiency in physical and visual generalization and its resilience under varying conditions. However, there is significant room for improvement in motion and semantic generalization. For instance, when presented with instructions such as ”Cut the banana” or ”Place the tomato on the plate,” the module exhibited inconsistent behavior. Examples of these scenarios are illustrated in Fig.  4.